OpenStudy (amberlykhan):

Explain how to find all the possible positive values of b such that x^2 + bx + 6 can be factored into binomial factors. Write the possible trinomials.

jimthompson5910 (jim_thompson5910):

did you mean this `x^2 + bx + 6` ?

OpenStudy (amberlykhan):

Yeah

jimthompson5910 (jim_thompson5910):

ok what are all the possible ways to multiply 2 numbers to get 6? here is one way: 2*3 = 6 what are some others?

OpenStudy (amberlykhan):

1 * 6

OpenStudy (amberlykhan):

-2 * -3 -1 * -6

jimthompson5910 (jim_thompson5910):

good, each pair of factors will add up to b so if we go with the pair 2 and 3, then b =2+3 = 5 that means x^2 + 5x + 6 is one possibility

OpenStudy (amberlykhan):

ok i understand that but what are binomial factors i got confused there

OpenStudy (amberlykhan):

wait

jimthompson5910 (jim_thompson5910):

binomial means 2 terms so anything where you have 2 terms added or subtracted example: x+5, or 2x-7

OpenStudy (amberlykhan):

is it like (x +2) (x +3)

jimthompson5910 (jim_thompson5910):

x^2 + 5x + 6 factors into (x+2)(x+3) notice how the 2 and 3 are there x+2 is one binomial factor x+3 is the other

jimthompson5910 (jim_thompson5910):

yes correct, that's one factorization

OpenStudy (amberlykhan):

sorry my computer was lagging

OpenStudy (amberlykhan):

thanks for clearing that up for me!